Two rectangular prisms are constructed. One measures 4cm times 6cm times x cm and the other measures 3cm times 8cm times y cm where both x and y are integers. If they have equal surface area, what is the smallest possible value of x+y?
Answer by mananth(16946)


Two rectangular prisms are constructed. One measures 4cm times 6cm times x cm and the other measures 3cm times 8cm times y cm where both x and y are integers. If they have equal surface area, what is the smallest possible value of x+y?
Surface area 1 = S1 (4,6,x)
Surface area of second = s2 (3,8,y)
S1=s2
S = 2( lw +wh +lh)
2(4*6+6*x+4*x) = 2(3*8+8*y+3y)
24+10x=24+11y
10x=11y
x/y = 11/10
x+y =21